An athelete takes 2.0 s to reach his maximum speed of 18.0 km/h. What is the magnitude of his average acceleration?

Grade:11

3 Answers

Kevin Nash
9 years ago
Sol. Initial velocity u = 0 (∴ starts from rest) Final velocity v = 18 km/hr = 5 sec (i.e. max velocity) Time interval t = 2 sec. ∴ Acceleration = aave = (v-u )/t = 5/2 = 2.5 m/s2.
Dhiraj
4 years ago
This question is very simple.
 
You just need to find average acceleration. 
Formula:v-u/t
U=0
V=18×5/18=5 m/s
T=2s
a=5/2 =2.5 m/s2
Vikas TU
4 years ago
Initial velocity=u= 0 m/s
Final velocity =v=18km/hr=18x5/18=5 m/s
time= 2sec
Average acceleration =a=?
Formula to be used : a= v-u/t
a=5-0/2 =2.5 m/s²
∴ Magnitude of his average acceleration is 2.5m/s²